# Large Language Models for Travel Behavior Prediction

### 30 Nov 2023 v1

@@ 4.1 Data
The model is evaluated on the Swissmetro stated preference survey data set (Bierlaire et al., 2001). The survey aims to analyze the impact of travel modal innovation in transportation, represented by the Swissmetro, a revolutionary maglev underground system, against the usual transport modes represented by car and train. The data contains 1,004 individuals with 9,036 responses. Users are asked to select from three travel modes (train, car, and Swissmetro) given the corresponding travel attributes.
The LLM-based prediction model is compared against three benchmark models: 1) multinomial logit (MNL), 2) random forest (RF), and 3) neural networks (NNs). MNL is the canonical travel behavior prediction tool, while random forest and NNs are proven to perform well empirically (Wang et al., 2021). Since the sample sizes are small, we do not include more advanced deep-learning models as benchmarks.
Considering the cost of the OpenAI API, We further down-sample from the whole data set. We use a balanced sampling to randomly select 1,000 samples as the training data (only for benchmark models) and 200 samples as the testing data set. Balanced sampling means the proportions of each travel mode in the final data set are almost the same.
The features and data statistics (combined training and testing) are shown in Table 1.
Table 1: Training and testing data statistics
@@
Train travel time (min) 160.6 74.0 34 794
Train cost (Swiss franc) 89.1 65.8 10 480
Car travel time (min) 139.1 98.0 32 1560
Car cost (Swiss franc) 87.9 49.4 10 520
Swissmetro travel time (min) 84.2 49.7 8 750
Swissmetro cost (Swiss franc) 110.2 83.5 14 768
Regular train user (Yes = 1) 0.34 0.47 0 1
@@ 4.2 Parameter settings
The specific LLM employed in the experiment is GPT-3.5 (version gpt-3.5-turbo-1106), which is one of the most advanced and widely used LLMs with open APIs. The prompts used are the same as Figure 7. We set the temperature to 0 to avoid randomness in the output. The benchmark models are implemented using the python sklearn package with the default hyper-parameters.
@@ 4.3.1 Prediction performance
All models are evaluated based on prediction accuracy and weighted F1-score. Results are shown in Table 2. The LLM-based prediction has shown competitive prediction performance as many commonly-used supervised learning models, though we do not provide any training samples.
Table 2: Prediction performance comparisons
Models Accuracy F1-score
LLM 0.655 0.648
MNL 0.645 0.639
RF 0.650 0.646
NN 0.685 0.676
@@ 4.3.2 Interpretability
A unique advantage of LLM-based prediction is the interpretability of the results. Table 3 shows several example outputs from the LLM. In Case 1, LLM can correctly predict the travel mode choice (while MNL cannot) and provide reasonable explanations. In Case 2, though LLM makes the wrong prediction, the explanations are reasonable. It talks about the reason for choosing a Car due to its lowest travel cost, and its travel time is not too different from other modes. In Case 3, we observe LLM’s reasoning errors, stating that the absence of a train annual pass makes the person choose the train over Swissmetro. In Case 4, we see a common hallucination problem of LLMs (Bang et al., 2023). It claims that the person prefers time-saving over cost-saving without any related inputs.
@@
Input:
{Travel time: {Train: 106, Car: 90, Swissmetro: 34}, Travel cost: {Train: 72, Car: 70, Swissmetro: 78}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction:
Swissmetro (MNL prediction is Car)
Actual:
Swissmetro
Reason:
Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Swissmetro offers the lowest travel time, saving 62% compared to Car and 68% compared to Train. Although Car has the lowest travel cost, the significant time saving of Swissmetro makes it the most likely choice for the person.
@@
Input:
{Travel time: {Train: 114, Car: 96, Swissmetro: 55}, Travel cost: {Train: 82, Car: 80, Swissmetro: 105}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction:
Car (MNL prediction is Car)
Actual:
Train
Reason:
Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Considering the trade-off between travel time and cost, the person is likely to choose the Car as it has the lowest travel cost and the difference in travel time compared to Swissmetro is not significant enough to outweigh the cost savings.
@@
Input:
{Travel time: {Train: 124, Car: 96, Swissmetro: 59}, Travel cost: {Train: 83, Car: 80, Swissmetro: 133}}. The person is a regular Train user. He/She does not own the Train annual pass.
Prediction:
Train (MNL prediction is Train)
Actual:
Swissmetro
Reason:
The person is a regular Train user, and regular Train users prefer to use Train. Although Swissmetro has the lowest travel time, the person’s preference for Train and the absence of a Train annual pass make it more likely for the person to choose Train over Swissmetro. Additionally, Car has the lowest travel cost, but the person’s regular use of Train and preference for it make it more likely for the person to choose Train over Car despite the cost difference.
@@
Input:
{Travel time: {Train: 95, Car: 130, Swissmetro: 92}, Travel cost: {Train: 29, Car: 44, Swissmetro: 32}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction:
Swissmetro (MNL prediction is Train)
Actual:
Train
Reason:
Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Swissmetro has the lowest travel time, saving 29% compared to Car and 3% compared to Train. Although Train has the lowest travel cost, the person’s attributes suggest a preference for time saving, making Swissmetro the most likely choice.
@@ 5 Conclusion and future work
In this study, we show the potential of using LLM for the complex travel behavior prediction task without using training samples. The LLM-based prediction has competitive performance as many commonly used supervised learning models such as MNL, RF, and NNs. Besides, LLM can also output explanations for the inference, though there are still reasoning errors and hallucination problems.
This paper presents a preliminary result, there is a lot of ongoing and future work. 1) First, more experiments are needed, including using larger data sets, conducting replicated experiments to mitigate random errors, and replicating the study in other travel behavior tasks such as travel purpose prediction. 2) The proposed prompt is composed of many components. Ablation studies can be conducted to test the effectiveness of different components in the prompt design. 3) The current prompt is a zero-shot prompt. Future studies can include some representative samples in the training data set as a few-shot prompt for in-context learning. The few-shot prompt needs to guide the LLMs to learn to quantify the value of time (or other important concepts for prediction) using the limited samples. 4) More experiments can be conducted on testing the performance of different LLMs as well.

### 10 Mar 2026 v2

@@ 4.1 Data
Mode choice prediction is evaluated on the Swissmetro stated preference survey data set [Bierlaire et al., 2001]. The survey aims to analyze the impact of travel modal innovation in transportation, represented by the Swissmetro, a revolutionary maglev underground system, against the usual transport modes represented by car and train. The data contains 1,004 individuals with 9,036 responses. Users are asked to select from three travel modes (train, car, and Swissmetro) given the corresponding travel attributes. Trip purpose prediction is evaluated using the 2017 US National Household Travel Survey (NHTS), which tracks daily trip, socio-demographic features, and trip purposes. A total of 4 trip purposes are included: working, social, recreation, and others.
For both datasets, we use balanced sampling to randomly select 1,000 entries for the large training set, 10 for the small training set (training sets are for benchmark models only), and 200 for the testing dataset. Balanced sampling means we ensure the proportion of all labels (i.e., mode choice and trip purposes) are the same. Results are derived from the average of 5 random samples.
The features and data statistics (combined all training and testing datasets) are shown in Tables 1 and 2.
Table 1: Swissmetro training and testing data statistics
@@
Train travel time (min) 161.2 77.3 31 1049
Train cost (Swiss franc) 87.4 65.1 8 576
Car travel time (min) 138.7 96.8 32 1560
Car cost (Swiss franc) 86.9 46.9 8 520
Swissmetro travel time (min) 85.7 53.5 8 790
Swissmetro cost (Swiss franc) 108.3 82.7 11 768
Regular train user (Yes = 1) 0.35 0.48 0 1
@@
Table 2: NHTS training and testing data statistics
Features Mean Std Min Max
With child (Yes = 1) 0.39 0.49 0 1
Primary activity last week is work (Yes = 1) 0.89 0.31 0 1
Trip duration less than 10 min (Yes = 1) 0.39 0.48 0 1
Trip duration greater than 25 min (Yes = 1) 0.28 0.45 0 1
Travel mode is walk/bike (Yes = 1) 0.08 0.27 0 1
Travel mode is drive (Yes = 1) 0.90 0.31 0 1
Travel mode is taxi/rideshare (Yes = 1) 0.00 0.07 0 1
Travel mode is transit (Yes = 1) 0.01 0.12 0 1
Travel on weekends (Yes = 1) 0.24 0.43 0 1
Only one person in trip (Yes = 1) 0.59 0.49 0 1
4.2 Baseline models
The LLM-based prediction model is compared against three benchmark models: 1) multinomial logit (MNL), 2) random forest (RF), and 3) neural networks (NNs). MNL is the canonical travel behavior prediction tool, while random forest and NNs are proven to perform well empirically [Wang et al., 2024c]. Since the sample sizes are small, we do not include more advanced deep-learning models as benchmarks.
@@ 4.3 Parameter settings
The specific LLM employed in the experiment is GPT-3.5 (version gpt-3.5-turbo-1106), GPT-4 (version gpt-4-turbo-2024-04-09), Llama 3.1 8B, and Llama 3.1 70B, which are all widely used LLMs with open APIs. The final prompts used for the two case studies are attached in Appendix A. We set the temperature to 0 to avoid randomness in the output. The benchmark models are implemented using the python sklearn package with the default hyper-parameters.
@@ 4.4.1 Prediction performance
All models are evaluated based on prediction accuracy and weighted F1-score. Results are shown in Table 3. Training pipeline large and small represent supervised learning with 1000 and 10 as training sample size, respectively. While zero-shot for LLM means directly prompting without any training. We see that supervised learning models (such as MNL, RF, and NN) trained on large datasets achieve the best overall performance for both mode choice and trip purpose prediction, confirming the effectiveness of data-driven training when sufficient labeled data are available. However, their performance deteriorates substantially when the training dataset is small, with notable drops in both accuracy and weighted F1-score across all model types.
Table 3: Combined model prediction results
Model Training pipeline Mode choice Trip purpose
Accuracy F1 score Accuracy F1 Score
MNL large 0.605 ± 0.036 0.595 ± 0.034 0.418 ± 0.043 0.401 ± 0.043
RF large 0.612 ± 0.031 0.613 ± 0.026 0.459 ± 0.046 0.451 ± 0.047
NN large 0.635 ± 0.031 0.637 ± 0.028 0.463 ± 0.047 0.455 ± 0.049
MNL small 0.469 ± 0.075 0.457 ± 0.080 0.313 ± 0.048 0.294 ± 0.058
MNL + Embedding small 0.494 ± 0.061 0.486 ± 0.057 0.327 ± 0.025 0.299 ± 0.029
RF small 0.455 ± 0.068 0.441 ± 0.074 0.250 ± 0.000 0.100 ± 0.000
RF + Embedding small 0.509 ± 0.077 0.498 ± 0.081 0.311 ± 0.025 0.278 ± 0.039
NN small 0.462 ± 0.044 0.455 ± 0.046 0.268 ± 0.042 0.213 ± 0.023
NN + Embedding small 0.532 ± 0.038 0.529 ± 0.038 0.296 ± 0.027 0.275 ± 0.031
LLM-GPT-3.5 zero-shot 0.586 ± 0.052 0.572 ± 0.058 0.402 ± 0.030 0.359 ± 0.040
LLM-GPT-4 zero-shot 0.570 ± 0.054 0.569 ± 0.054 0.401 ± 0.014 0.381 ± 0.020
LLM-Llama-3.1 8B zero-shot 0.551 ± 0.047 0.539 ± 0.041 0.322 ± 0.008 0.243 ± 0.011
LLM-Llama-3.1 70B zero-shot 0.540 ± 0.038 0.532 ± 0.041 0.408 ± 0.022 0.343 ± 0.022
In contrast, LLM-based zero-shot predictions—without using any training samples—can outperform supervised models trained on limited data. This highlights the practical value of LLMs in real-world scenarios where labeled data are scarce or unavailable. Despite the absence of task-specific training, LLMs are able to leverage their pretrained knowledge to deliver competitive predictions.
Furthermore, when small training datasets are available, training supervised models with LLM-generated text embeddings consistently improves performance compared to using manually engineered features alone. This suggests that LLM embeddings capture richer semantic information that is difficult to encode through traditional feature design, thereby enhancing model generalization under data-constrained settings. Overall, these findings indicate that LLMs and their embeddings offer a promising alternative or complement to conventional supervised learning, particularly in low-data regimes.
@@ 4.4.2 Interpretability
Input: {Travel time: {Train: 106, Car: 90, Swissmetro: 34}, Travel cost: {Train: 72, Car: 70, Swissmetro: 78}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction: Swissmetro (MNL prediction is Car)
Actual: Swissmetro
Reason: Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Swissmetro offers the lowest travel time, saving 62% compared to Car and 68% compared to Train. Although Car has the lowest travel cost, the significant time saving of Swissmetro makes it the most likely choice for the person.
@@
Input: {Travel time: {Train: 114, Car: 96, Swissmetro: 55}, Travel cost: {Train: 82, Car: 80, Swissmetro: 105}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction: Car (MNL prediction is Car)
Actual: Train
Reason: Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Considering the trade-off between travel time and cost, the person is likely to choose the Car as it has the lowest travel cost and the difference in travel time compared to Swissmetro is not significant enough to outweigh the cost savings.
@@
Input: {Travel time: {Train: 124, Car: 96, Swissmetro: 59}, Travel cost: {Train: 83, Car: 80, Swissmetro: 133}}. The person is a regular Train user. He/She does not own the Train annual pass.
Prediction: Train (MNL prediction is Train)
Actual: Swissmetro
Reason: The person is a regular Train user, and regular Train users prefer to use Train. Although Swissmetro has the lowest travel time, the person’s preference for Train and the absence of a Train annual pass make it more likely for the person to choose Train over Swissmetro. Additionally, Car has the lowest travel cost, but the person’s regular use of Train and preference for it make it more likely for the person to choose Train over Car despite the cost difference.
@@
Input: {Travel time: {Train: 95, Car: 130, Swissmetro: 92}, Travel cost: {Train: 29, Car: 44, Swissmetro: 32}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction: Swissmetro (MNL prediction is Train)
Actual: Train
Reason: Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Swissmetro has the lowest travel time, saving 29% compared to Car and 3% compared to Train. Although Train has the lowest travel cost, the person’s attributes suggest a preference for time saving, making Swissmetro the most likely choice.
A unique advantage of LLM-based prediction is the interpretability of the results. Table 4 shows several example outputs from the LLM. In Case 1, LLM can correctly predict the travel mode choice (while MNL cannot) and provide reasonable explanations. In Case 2, though LLM makes the wrong prediction, the explanations are reasonable. It talks about the reason for choosing a Car due to its lowest travel cost, and its travel time is not too different from other modes. In Case 3, we observe LLM’s reasoning errors, stating that the absence of a train annual pass makes the person choose the train over Swissmetro. In Case 4, we see a common hallucination problem of LLMs [Bang et al., 2023]. It claims that the person prefers time-saving over cost-saving without any related inputs.
4.4.3 Ablation studies
In this section, we show how the exclusion of different components in the prompt would affect the performance. For simplicity, only the ablation of mode choice cases is tested with two GPT models. Note that, the ablation study still keeps all the information (same as the information given to supervised learning model), but with different prompt design. The meaning of different test cases are as follows:
  • 1.
    Exclude guide of thinking: We remove the sentences of “Swissmetro has the lowest travel time. Choosing it will save 39% travel time compared to …” from the prompt, which is used to guide the LLM to compare cost and travel time.
  • 2.
    Exclude domain knowledge: We remove the sentences of “Please consider the following aspects: 1. People are more likely to choose …” from the prompt, which provides domain knowledge on how people typically make their decisions.
  • 3.
    Change structural data to non-structural data: Instead of using structural data as inputs (for example, {Travel time: {Train: 202, Car: 160, Swissmetro: 97}}), we directly tell the LLM about this information with descriptive sentences (e.g., the travel time of train is 202 minutes).
Table 5: Ablation study results for the Swissmetro model choice dataset
Model Prompt design Performance
Accuracy (Decrease %)1 F1 Score
LLM-GPT-3.5 Full prompt 0.586 ± 0.052 0.572 ± 0.058
Exclude guide of thinking 0.552 ± 0.036 (-5.8%) 0.526 ± 0.039
Exclude domain knowledge 0.583 ± 0.052 (-0.5%) 0.571 ± 0.052
Change structural data to non-structural data 0.560 ± 0.037 (-4.4%) 0.538 ± 0.040
LLM-GPT-4 Full prompt 0.570 ± 0.054 0.569 ± 0.054
Exclude guide of thinking 0.565 ± 0.044 (-0.8%) 0.562 ± 0.043
Exclude domain knowledge 0.569 ± 0.050 (-0.2%) 0.568 ± 0.052
Change structural data to non-structural data 0.567 ± 0.044 (-0.5%) 0.565 ± 0.043
1: Values in the parenthesis represents the mean accuracy decrease compared to the full prompt.
As shown in Table 5, the ablation results indicate that prompt design plays a markedly different role for LLM-GPT-3.5 and LLM-GPT-4. For LLM-GPT-3.5, excluding the guide of thinking leads to the largest performance degradation, with accuracy dropping by 5.8%, suggesting that explicit comparative reasoning cues are critical for this model to effectively weigh attributes such as travel time and cost. Changing structural data to non-structural, descriptive inputs also results in a performance decrease (4.4%), highlighting the importance of well-organized, structured information for LLM-GPT-3.5. In contrast, excluding domain knowledge has a negligible impact (0.5% decrease), which may indicate that such decision-making heuristics are already embedded in the model’s general reasoning capability. For LLM-GPT-4, performance remains largely stable across all ablation settings, with accuracy decreases below 1% in all cases. This robustness suggests that GPT-4 is less sensitive to prompt engineering details and can effectively extract and utilize relevant information regardless of explicit reasoning guidance, domain knowledge cues, or data representation format, reflecting its more advanced reasoning and information integration capabilities.
@@ 5 Conclusion and future work
This study investigates the potential of large language models (LLMs) as an alternative and complement to conventional travel behavior prediction methods. We propose and evaluate two LLM-based frameworks: a zero-shot prompting approach that directly generates predictions without task-specific training data, and a hybrid approach that integrates LLM-generated text embeddings with classical supervised learning models. Empirical results on mode choice and trip purpose prediction demonstrate that while traditional supervised models such as MNL, random forest, and neural networks achieve the best performance when large labeled datasets are available, their effectiveness degrades substantially in low-data regimes. In contrast, LLM-based zero-shot predictions achieve competitive performance without any training data and consistently outperform supervised models trained on very small samples. Moreover, incorporating LLM embeddings into supervised models significantly improves prediction accuracy and F1 scores under data-scarce conditions, indicating that LLMs capture high-level semantic information that is difficult to encode through manual feature engineering.
Beyond predictive performance, LLMs offer a distinct advantage in interpretability by providing natural language explanations for their predictions. Our qualitative analysis shows that these explanations are often reasonable and aligned with travel behavior theory, even when the predictions are incorrect. However, we also observe cases of reasoning errors and hallucinations, underscoring the need for caution when using LLM-generated explanations in decision-critical applications. The ablation study further highlights the importance of prompt design, particularly for less advanced models such as GPT-3.5, where explicit reasoning guidance and structured inputs play a significant role in prediction quality. In contrast, GPT-4 exhibits strong robustness to prompt variations, suggesting improved internal reasoning and information integration capabilities.
This work represents an initial exploration of LLMs for travel behavior prediction, and several directions remain for future research. First, more extensive experiments are needed using larger and more diverse datasets, including repeated trials to reduce randomness and evaluations across additional travel behavior tasks, such as departure time choice, route choice, and activity scheduling. Second, while this study includes an ablation analysis of prompt components, more systematic investigations into prompt optimization and standardization are needed to better understand how different forms of reasoning guidance, domain knowledge, and data representation affect LLM performance. Third, future work can explore few-shot and in-context learning strategies, where carefully selected representative samples are provided to guide LLMs in learning quantitative trade-offs such as the value of travel time or cost sensitivity from limited data. Finally, broader comparisons across a wider range of open-source and proprietary LLMs, as well as investigations into methods for reducing hallucinations and improving explanation faithfulness, would further clarify the practical applicability of LLMs in transportation modeling.
[…]
Appendix A Prompts
The final prompts for the travel mode prediction and trip purpose predictions are shown in Table 6. Both are presented with an example. All changeable parts are underscored, where different samples will have different values.
Table 6: Final prompts used for predictions
Prompt for travel mode prediction
Your task is to predict a person’s travel mode choice from Train, Car, and Swissmetro based on travel time, travel cost information of each mode, and the person’s attributes. The travel time and cost for each mode is expressed as the following dictionary format: {Travel time: {Train: 94, Car: 56, Swissmetro: 36}, Travel cost: {Train: 62, Car: 36, Swissmetro: 71}} Swissmetro has the lowest travel time, Choosing it will save 36% travel time compared to Car and save 62% travel time compared to Train. Car has the lowest travel cost. Choosing it will save 49% travel cost compared to Swissmetro and save 42% travel cost compared to Train. The person is a regular Train user. He/She does not own the Train annual pass. Please infer what is the mostly likely travel mode that the person will choose. Organize your answer in a JSON object with two keys: ”prediction” (the predicted travel mode) and ”reason” (explanation that supports your inference). Please consider the following aspects: 1. People are more likely to choose a travel mode with less travel cost and travel time, especially those with significant cost or time saving. The trade off between time and cost can be quantified using value of time. 2. Regular Train users prefer to use Train. 3. Owners of Train annual pass are more likely to choose Train.
Prompt for trip purpose prediction
Your task is to infer the purpose of a person’s trip, classifying it into one of four categories: “working”, “shopping”, “social”, and “others”. You will make the inference based on the following information related to the socio-demographic information and trip information of the traveler. The person lives in a household with child. As of last week, His/her primary activity is working. The trip happens on weekdays and starts after 5pm. The trip lasts less than 10 minutes. The mode of transportation is not walking or biking. There were 2 or more people on the trip. Please also consider the following reasoning logics to support your inference: 1. Trips that happen on weekends are more likely to be social or shopping trips 2. Trips that happen before 10am are more likely to be working trips. Trips that happen after 5pm are more likely to be social trips 3. A trip with only one person is highly likely to be a working trip 4. Trips with walking or biking transportation mode are highly likely to be social trips 5. Working and social trip duration is usually longer than 25 min 6. People with child are more likely to trip purpose of ”others”, such as school or daycare 7. If the person’s primary activity last week is working, it is more likely to be a working trip Please follow the above logics strictly, especially not violate the rules. Infer what is the mostly likely trip purpose. Organize your final answer in a JSON object with two keys: ”prediction” (the predicted travel purpose) and ”reason” (explanation that supports your inference in one paragraph).